### v1

@@ 4.1. Research Questions
  • RQ3: How is the effectiveness of entanglement guidance for quantum adversarial testing?
To demonstrate the efficacy of QuanTest in capturing erroneous behaviors of QNN systems, we conducted tests using QuanTest on 9 representative mainstream QNN models. Subsequently, an analysis was performed on the generation of test samples by QuanTest across varying numbers of iterations. Finally, we evaluated the performance of QuanTest in quantum systems with varying dimensions. To answer RQ2, we analyze the fidelity measure and trace distance of the test samples generated by QuanTest, without imposing any similarity constraints. To address RQ3, we conducted control experiments pertaining to entanglement-guided settings and analyzed the performance of QuanTest under the guidance of QEA with different values of $k$.
@@ 4.2.3. Data Encoding
Amplitude encoding maps the feature information of classical data to the amplitudes of quantum states. For the input classical data ${\bf{x}}\in\mathbb{R}^{N}$, its vector elements are encoded and mapped to the amplitudes of $\lceil\log_{2}N\rceil$ qubits: $|{\bf{x}}\rangle=\sum_{i=1}^{N}x_{i}|i\rangle$, where $\{|i\rangle\}$ represents a set of computational basis states in the Hilbert space. Since the feature information of classical data is represented as the amplitudes of a quantum state, the classical input must satisfy the normalization condition: $\|{\bf{x}}\|^{2}=\sum_{i}|x_{i}|^{2}=1$. We require the use of 256 amplitudes of the 8-qubit system to encode this data information, as each image in the MNIST and Fashion-MNIST datasets is two dimensional and contains 16$\times$16 pixels. Note that QuntTest also possesses universality and extensibility for other encoding strategies.
@@ 4.2.4. Test Samples Generation
We use adversarial attack strategies FGSM (Goodfellow et al., 2015) and DLFuzz (Guo et al., 2018) under the premise of an entanglement-guided approach to search for erroneous behaviors and generate quantum adversarial examples, for testing QNN models on the MNIST and Fashion-MNIST datasets encoded as quantum states. We have adapted these adversarial attack strategies with quantum-based adjustments to enable them to process quantum state inputs and the information extracted within QNN systems. Essentially, both of these strategies rely on a gradient-based manner to solve the optimization problem. These adversarial examples have imperceptible perturbations that are difficult for humans to detect, and demonstrate considerable effectiveness in revealing robustness issues in QNN systems.
In addition, we have prepared random noise samples generated by applying random coherent noise to original samples for subsequent comparative evaluation. We realized coherent noise by constructing a universal quantum rotation gate with Gaussian noise perturbation on the rotation parameters. This allows us to apply it as a random perturbation operator $U_{\Delta}(\sigma)$ on quantum original samples in the form of a quantum circuit, where $\sigma$ represents the strength of the noise, reflecting the magnitude of the standard deviation in the normal distribution.
@@ 4.3. Capabilities of Erroneous Behavior Detection
Table 2 shows the overall results obtained with QuanTest and random coherent noise, where Gen Rate represents the test sample generation rate. The best performance is typeset in bold. Test samples generated by each strategy exhibit comparable similarities, as restrictions are imposed on the similarity metrics. The maximum difference for AFM does not exceed 3.89%, and for ATD, it does not exceed 0.14. This allows the evaluation of the test sample generation rates for each strategy with minimal interference. It can be observed that the test sample generation rate of coherent noise is generally low, reaching a maximum of only 4.58%. In contrast, QuanTest achieves a maximum test sample generation rate of 84.59% using the FGSM attack strategy, while the use of the DLFuzz attack strategy can result in a success rate of up to 96.87%. The test sample generation rate of the DLFuzz strategy is 9.53% to 26.43% higher than that of the FGSM strategy across various models.
@@
Subsequently, we calculated the test sample generation rate under different iteration counts using the entanglement-guided DLFuzz strategy when no similarity constraints were imposed (i.e., Gen Rate=100%), as shown in Figure 6. The results indicate that QuanTest based on the DLFuzz strategy is capable of generating the majority of adversarial examples after 5-10 iterations. In contrast, QuanTest based on the FGSM strategy requires slightly more iterations. This may be related to the effectiveness of the adversarial attack strategy. In addition, the speed of test sample generation for QCNN is generally faster than other models, which may be related to the quality of the model construction.
We also investigate the performance of QuanTest in capturing erroneous behaviors in quantum systems of different dimensions. We used 784 amplitudes from the 10-qubit system to prepare quantum states for images of size 28$\times$28 pixels in the dataset, with the last 240 bits of the state vector padded with zeros to ensure a length of $2^{10}$. We extended the pre-trained QNN model accordingly, increasing the number of quantum gates and trainable parameters by 0.25 times due to the increase in the number of qubits. The training was terminated when the model achieved performance similar to the QNN model corresponding to 16$\times$16 pixel images in the classification. This was done to ensure fairness in the comparative experiments. For each classification category in the classification task, we randomly selected 500 samples of size 16$\times$16 pixels and 28$\times$28 pixels. fidelity measure ¿ 95% or trace distance ¡ 0.30 were chosen as the criteria for adversarial example detection.
@@ 4.4. Quality of Generated adversarial examples
To address the RQ2, we conducted a statistical analysis of the fidelity measure and trace distance between the quantum adversarial examples generated by QuanTest using different adversarial attack strategies and the quantum original samples. QuanTest generated 600 adversarial examples for each classification task without imposing any similarity constraints. In other words, we assessed the quality of the adversarial examples generated by QuanTest under the condition of Gen Rate=100%. This condition setting enables our statistics to cover a wide range of scenarios, including extreme cases, thereby enhancing the comprehensiveness of the evaluation.
@@
Answer to RQ2: The adversarial examples generated by QuanTest exhibit high quality and perform well in terms of both fidelity measure and trace distance.
@@ 4.5. Effectiveness of Entanglement Guidance
Table 4. Performance of the adversarial testing with and without entanglement guidance.
@@
QEA(%) ($k=1$)
Gen Rate(%)
AFM(%)
ATD
Fashion[1,4] CCQC
$\mathcal{D}$
49.55
71.49
94.84 0.3061
$\mathcal{D}+\mathcal{G}$
57.12 91.17
93.42
0.3452
$\mathcal{F}$
51.39
11.78
99.00 0.1305
$\mathcal{F}+\mathcal{G}$
57.88 19.60
97.41
0.2117
QCL
$\mathcal{D}$
54.10
70.00
96.15 0.2587
$\mathcal{D}+\mathcal{G}$
61.45 85.86
94.33
0.3212
$\mathcal{F}$
55.44
10.88
99.21 0.1170
$\mathcal{F}+\mathcal{G}$
62.12 13.69
98.04
0.1834
QCNN
$\mathcal{D}$
38.08
75.31
97.07 0.2225
$\mathcal{D}+\mathcal{G}$
44.84 89.22
95.28
0.2890
$\mathcal{F}$
38.19
21.25
99.06 0.1272
$\mathcal{F}+\mathcal{G}$
44.93 36.08
97.61
0.2049
MNIST[0,1,2] CCQC
$\mathcal{D}$
16.25
33.11
98.98 0.1247
$\mathcal{D}+\mathcal{G}$
24.53 41.90
97.32
0.2093
$\mathcal{F}$
17.05
14.22
99.70 0.6983
$\mathcal{F}+\mathcal{G}$
24.66 15.55
98.78
0.1401
QCL
$\mathcal{D}$
19.93
48.60
98.85 0.1342
$\mathcal{D}+\mathcal{G}$
28.00 54.33
97.37
0.2070
$\mathcal{F}$
19.80
18.28
99.67 0.739
$\mathcal{F}+\mathcal{G}$
26.82 21.39
98.81
0.1384
QCNN
$\mathcal{D}$
18.09
69.34
98.66 0.1480
$\mathcal{D}+\mathcal{G}$
22.89 71.14
98.00
0.1794
$\mathcal{F}$
18.19
36.08
99.60 0.8282
$\mathcal{F}+\mathcal{G}$
23.34 42.07
98.88
0.1364
MNIST[3,6] CCQC
$\mathcal{D}$
19.93
55.33
96.92 0.2312
$\mathcal{D}+\mathcal{G}$
28.07 76.65
95.10
0.2958
$\mathcal{F}$
17.10
12.36
99.12 0.1263
$\mathcal{F}+\mathcal{G}$
26.10 21.29
97.78
0.1973
QCL
$\mathcal{D}$
21.32
60.37
98.34 0.1676
$\mathcal{D}+\mathcal{G}$
31.79 61.80
96.98
0.2314
$\mathcal{F}$
21.55
14.51
99.59 0.8431
$\mathcal{F}+\mathcal{G}$
31.55 15.52
98.59
0.1553
QCNN
$\mathcal{D}$
16.05
83.92
97.16 0.2213
$\mathcal{D}+\mathcal{G}$
23.67 89.37
95.67
0.2773
$\mathcal{F}$
15.34
18.39
99.16 0.1212
$\mathcal{F}+\mathcal{G}$
23.34 28.63
98.06
0.1837

### v2

@@ 4.1. Research Questions
  • RQ3: How is the effectiveness of entanglement guidance for QuanTest?
  • RQ4: How adversarial examples generated by QuanTest affect the sampling cost of the QNN system?
  • RQ5: How does QuanTest perform compared to black-box baseline?
To demonstrate the efficacy of QuanTest in capturing erroneous behaviors of QNN systems, we conducted tests using QuanTest on 9 representative mainstream QNN models. Subsequently, an analysis was performed on the generation of test samples by QuanTest across varying numbers of iterations. Finally, we evaluated the performance of QuanTest in quantum systems with varying dimensions. To answer RQ2, we analyze the fidelity measure and trace distance of the test samples generated by QuanTest, without imposing any similarity constraints. Additionally, we explored the performance of the test samples in improving model performance after retraining. To address RQ3, we conducted control experiments pertaining to entanglement-guided settings and analyzed the performance of QuanTest under the guidance of QEA with different values of $k$. For RQ4, we model the measurement of the predicted probability qubit as a sampling problem in a Bernoulli distribution, based on which we explored the impact of the adversarial examples generated by QuanTest on the QNN’s prediction under different sampling times. Finally, we compare the performance of QuanTest with state-of-the-art black-box baseline, and discuss the usage scenarios of QuanTest.
@@ 4.2.3. Data Encoding
Amplitude encoding maps the feature information of classical data to the amplitudes of quantum states. For the input classical data ${\bf{x}}\in\mathbb{R}^{N}$, its vector elements are encoded and mapped to the amplitudes of $\lceil\log_{2}N\rceil$ qubits: $|{\bf{x}}\rangle=\sum_{i=1}^{N}x_{i}|i\rangle$, where $\{|i\rangle\}$ represents a set of computational basis states in the Hilbert space. Since the feature information of classical data is represented as the amplitudes of a quantum state, the classical input must satisfy the normalization condition: $\|{\bf{x}}\|^{2}=\sum_{i}|x_{i}|^{2}=1$.
We require the use of 256 amplitudes of the 8-qubit system to encode this data information, as each image in the MNIST and Fashion-MNIST datasets is two dimensional and contains 16$\times$16 pixels. Note that QuanTest also possesses universality and extensibility for other encoding strategies.
@@ 4.2.4. Test Samples Generation
We use adversarial attack strategies FGSM (Goodfellow et al., 2015) and DLFuzz (Guo et al., 2018) under the premise of an entanglement-guided approach to search for erroneous behaviors and generate quantum adversarial examples, for testing QNN models on the MNIST and Fashion-MNIST datasets encoded as quantum states. We have adapted these adversarial attack strategies with quantum-based adjustments to enable them to process quantum state inputs and the information extracted within QNN systems. Essentially, both of these strategies rely on a gradient-based manner to solve the optimization problem. These adversarial examples have imperceptible perturbations that are difficult to detect, and demonstrate considerable effectiveness in revealing robustness issues in QNN systems.
In addition, we have prepared random noise samples generated by applying random coherent noise to original samples for subsequent comparative evaluation. We realized coherent noise by constructing a universal quantum rotation gate with Gaussian noise perturbation on the rotation parameters. This allows us to apply it as a random perturbation operator $U_{\Delta}(\sigma)$ on quantum original samples in the form of a quantum circuit, where $\sigma$ represents the strength of the noise, reflecting the magnitude of the standard deviation in the normal distribution. Note that all samples need to be able to flip the model’s predictions and pass quality evaluation to be output as test samples.
4.2.5. Evaluation Metrics
To assess the effectiveness of QuanTest, we define the test sample generation rate (Gen_Rate), which represents the ratio of the original inputs the method has successfully produced quantum test samples:
[…]
where $S_{adv}$ represents the generated adversarial examples that have successfully passed quality evaluation (Algorithm 1 line 1, 15-18), and $inputs$ represents the original input set (Algorithm 1 line 2). For the quality of the generated test samples, we introduce the AFM and ATD defined in Section 4.2.3 to assist us in evaluation.
@@ 4.3. Capabilities of Erroneous Behavior Detection
Table 2 shows the overall results obtained with QuanTest and random coherent noise, where Gen_Rate represents the test sample generation rate. The best performance is typeset in bold. Test samples generated by each strategy exhibit comparable similarities, as restrictions are imposed on the similarity metrics. The maximum difference for AFM does not exceed 3.89%, and for ATD, it does not exceed 0.14. This allows the evaluation of the test sample generation rates for each strategy with minimal interference. It can be observed that the test sample generation rate of coherent noise is generally low, reaching a maximum of only 4.58%. In contrast, QuanTest achieves a maximum test sample generation rate of 84.59% using the FGSM attack strategy, while the use of the DLFuzz attack strategy can result in a success rate of up to 96.87%. The test sample generation rate of the DLFuzz strategy is 9.53% to 26.43% higher than that of the FGSM strategy across various models.
@@
Subsequently, we calculated the test sample generation rate under different iteration counts using the entanglement-guided DLFuzz strategy when no similarity constraints were imposed (i.e., Gen_Rate=100%), as shown in Figure 6. The results indicate that QuanTest based on the DLFuzz strategy is capable of generating the majority of adversarial examples after 5-10 iterations. In contrast, QuanTest based on the FGSM strategy requires slightly more iterations. This may be related to the effectiveness of the adversarial attack strategy. In addition, the speed of test sample generation for QCNN is generally faster than other models, which may be related to the quality of the model construction.
We also investigate the performance of QuanTest in capturing erroneous behaviors when scaling to higher-dimensional quantum systems. We used 784 amplitudes from the 10-qubit system to prepare quantum states for images of size 28$\times$28 pixels in the dataset, with the last 240 bits of the state vector padded with zeros to ensure a length of $2^{10}$. We extended the pre-trained QNN model accordingly, increasing the number of quantum gates and trainable parameters by 0.25 times due to the increase in the number of qubits. The training was terminated when the model achieved performance similar to the QNN model corresponding to 16$\times$16 pixel images in the classification. This was done to ensure fairness in the comparative experiments. For each classification category in the classification task, we randomly selected 500 samples of size 16$\times$16 pixels and 28$\times$28 pixels. fidelity measure ¿ 95% or trace distance ¡ 0.30 were chosen as the criteria for adversarial example detection.
@@ 4.4. Quality of Generated Adversarial Examples
The quality of test samples can be multifaceted. To address the RQ2, we measure the quality of adversarial examples generated by QuanTest by evaluating the similarity between adversarial examples and original samples, as well as their improvement in enhancing the performance of QNN systems after retraining. We conducted a statistical analysis of the fidelity measure and trace distance between the quantum adversarial examples generated by QuanTest using different adversarial attack strategies and the quantum original samples. QuanTest generated 600 adversarial examples for each classification task without imposing any similarity constraints. In other words, we assessed the quality of the adversarial examples generated by QuanTest under the condition of Gen_Rate=100%. This condition setting enables our statistics to cover a wide range of scenarios, including extreme cases, thereby enhancing the comprehensiveness of the evaluation.
@@
We subsequently merged the generated test samples from Section 4.3 into the corresponding training sets of classification tasks, and retrained the QNN models with the same settings to investigate whether the test samples generated by QuanTest can be used for retraining to further improve model performance. In addition to using test samples generated by single strategies to constitute enhanced datasets, we also merged the test samples generated by QuanTest based on two adversarial attack strategies, DLFuzz and FGSM, with the original dataset to form a new enhanced dataset, aiming to explore the performance of mixed test samples generated by multiple strategies.
Figure 8 shows the accuracy of the model after retraining on different enhanced datasets. Here the data are averaged over three times of the same retraining process to mitigate the impact arising from uncertainty. It can be seen that the test samples generated by QuanTest can enhance the training effect of the model by adding more corner cases. Models with lower original accuracy often gain more improvement after retraining, especially the enhanced training set composed of mixed test samples, which can bring up to a maximum of 7.1% improvement. This means that adversarial examples generated based on different attack strategies can discover different types of erroneous behaviors, thereby improving the robustness of QNN systems.
Figure 8. Accuracy of three QNN models for the MNIST binary classification task before and after retraining.
[…]
The accuracy of three QNN models for the MNIST binary classification task before and after retraining.
Answer to RQ2: The adversarial examples generated by QuanTest exhibit high quality and perform well in terms of both fidelity measure and trace distance. They can be utilized for retraining to enhance the robustness of QNN systems.
@@ 4.5. Effectiveness of Entanglement Guidance
Table 4. Performance of QuanTest with and without entanglement guidance.
@@
QEA(%) ($k=1$) Gen_Rate(%) AFM(%) ATD
Fashion[1,4] CCQC $\mathcal{D}$ 49.55 71.49 94.84 0.3061
$\mathcal{D}+\mathcal{G}$ 57.12 91.17 93.42 0.3452
$\mathcal{F}$ 51.39 11.78 99.00 0.1305
$\mathcal{F}+\mathcal{G}$ 57.88 19.60 97.41 0.2117
QCL $\mathcal{D}$ 54.10 70.00 96.15 0.2587
$\mathcal{D}+\mathcal{G}$ 61.45 85.86 94.33 0.3212
$\mathcal{F}$ 55.44 10.88 99.21 0.1170
$\mathcal{F}+\mathcal{G}$ 62.12 13.69 98.04 0.1834
QCNN $\mathcal{D}$ 38.08 75.31 97.07 0.2225
$\mathcal{D}+\mathcal{G}$ 44.84 89.22 95.28 0.2890
$\mathcal{F}$ 38.19 21.25 99.06 0.1272
$\mathcal{F}+\mathcal{G}$ 44.93 36.08 97.61 0.2049
MNIST[0,1,2] CCQC $\mathcal{D}$ 16.25 33.11 98.98 0.1247
$\mathcal{D}+\mathcal{G}$ 24.53 41.90 97.32 0.2093
$\mathcal{F}$ 17.05 14.22 99.70 0.6983
$\mathcal{F}+\mathcal{G}$ 24.66 15.55 98.78 0.1401
QCL $\mathcal{D}$ 19.93 48.60 98.85 0.1342
$\mathcal{D}+\mathcal{G}$ 28.00 54.33 97.37 0.2070
$\mathcal{F}$ 19.80 18.28 99.67 0.739
$\mathcal{F}+\mathcal{G}$ 26.82 21.39 98.81 0.1384
QCNN $\mathcal{D}$ 18.09 69.34 98.66 0.1480
$\mathcal{D}+\mathcal{G}$ 22.89 71.14 98.00 0.1794
$\mathcal{F}$ 18.19 36.08 99.60 0.8282
$\mathcal{F}+\mathcal{G}$ 23.34 42.07 98.88 0.1364
MNIST[3,6] CCQC $\mathcal{D}$ 19.93 55.33 96.92 0.2312
$\mathcal{D}+\mathcal{G}$ 28.07 76.65 95.10 0.2958
$\mathcal{F}$ 17.10 12.36 99.12 0.1263
$\mathcal{F}+\mathcal{G}$ 26.10 21.29 97.78 0.1973
QCL $\mathcal{D}$ 21.32 60.37 98.34 0.1676
$\mathcal{D}+\mathcal{G}$ 31.79 61.80 96.98 0.2314
$\mathcal{F}$ 21.55 14.51 99.59 0.8431
$\mathcal{F}+\mathcal{G}$ 31.55 15.52 98.59 0.1553
QCNN $\mathcal{D}$ 16.05 83.92 97.16 0.2213
$\mathcal{D}+\mathcal{G}$ 23.67 89.37 95.67 0.2773
$\mathcal{F}$ 15.34 18.39 99.16 0.1212
$\mathcal{F}+\mathcal{G}$ 23.34 28.63 98.06 0.1837
@@
5. Discussion
5.1. Effect of QuanTest on Sampling Cost
In this section, we explore how adversarial examples generated by QuanTest affect the sampling cost of QNN system. In order to read out the state of a qubit, we must measure it. Unlike quantum state tomography, which reconstructs the complete quantum state through measurements, obtaining predictive estimates from the measurement results output by QNN models is a simpler problem. Taking binary classification as an example, at least one qubit is needed to encode the prediction results of QNN model $\mathcal{M}$. Assuming that the qubit is in state $|\mathcal{M}(x)\rangle=|1\rangle$ the prediction is 1, and if it is in the state $|\mathcal{M}(x)\rangle=|0\rangle$ the prediction is 0. Superposition can be interpreted as providing probability outputs regarding the uncertainty of the result. At this point, $\mathcal{M}(x)=p$ represents the probability that the model’s prediction of $y=1$ for input $x$. In the case of error-free quantum algorithms, measuring a single predicted probability qubit of the QNN model $\mathcal{M}$ can evolve into a problem of sampling from a Bernoulli distribution.
The most straightforward and intuitive method to obtain an estimate $\hat{p}$ for $p$ in numerical simulations is maximum likelihood estimation, with the result $\bar{p}$ being the average of all sampled results. However, in real scenarios, what concerns us more is how many times $N$ we need to sample from the single qubit measurements in order to estimate the confidence interval $[\hat{p}-\epsilon,\hat{p}+\epsilon]$ for $p$ with a confidence level of $1-\alpha$. A refined estimation approach is the Wilson score interval, and the error $\epsilon$ can be calculated as:
(13) $\epsilon=\frac{z_{\alpha/2}}{1+\frac{z^{2}_{\alpha/2}}{N}}(\frac{\bar{p}(1-\bar{p})}{N}+\frac{z^{2}_{\alpha/2}}{4N^{2}})^{\frac{1}{2}},$
where $z_{\alpha/2}$ is the quantile (right-tail probabilities) of the stand normal distribution. With a confidence level of 99% (i.e., $z_{\alpha/2}\approx 2.58$), the relationship between the sampling cost and $\bar{p}$ can be obtained as shown in Figure 10 (a). It can be seen that the sampling cost reaches its maximum value when $p=0.5$.
Figure 10. Effect of sampling cost on QuanTest. (a) Relationship between the sampling cost and the mean value $\bar{p}$ for different errors $\epsilon$. (b) Error rate of seed sample predictions by the QNN system under different sampling costs.
[…]
The test samples generated by QuanTest can detect erroneous behavior in the QNN system. When the adversarial examples cause changes in the predicted labels of the tested QNN model, QuanTest will cease to apply perturbations to the samples and output the samples that have passed the quality evaluation as test samples (Algorithm 1 lines 15-18). At this point, the tested QNN model’s probability estimate of predicting $y=1$ for the test samples is close to 0.5. This means that the sampling cost required by the tested QNN system to predict the test samples generated by QuanTest with high confidence will increase significantly, especially in the face of the real system with fixed and limited sampling times, the test samples are more likely to trigger the system to make incorrect predictions, allowing QuanTest to test the robustness of QNN systems at multiple stages. the generated samples are more likely to cause the system to misestimate the prediction results.
To verify this, we first mark as seed samples 500 intermediate samples that were iterated 5 times in QuanTest but still failed to flip the model predictions, and then explore the influence of these seed samples on the prediction accuracy of QNN systems under different sampling times. We chose to evaluate the ternary classification task of MNIST with a low test sample generation rate to make it easier for us to obtain sufficient seed samples. Note that these seed samples cannot flip the prediction of the model in the case of ideal expectations (i.e., the number of samples tends to be $\infty$). Figure 10 (b) shows the error rate of three QNN models as the number of samples varies from $10^{1}$ to $10^{5}$, we repeated the classification for each model 10 times to mitigate randomness, where solid lines represent the average results and shadows represent the standard deviation. The results confirm our modeling-based inference that when the number of samples is small, the QNN system is more likely to exhibit erroneous behavior, which may also lead to QuanTest generating more test samples than expected.
To further assess the false negatives and false positives introduced by reducing the number of samples, we additionally explored the differences in the effect of QuanTest on the quality indicators of the QNN system under different levels of sampling costs. The classification task is set to the ternary classification on MNIST, and the number of iterations for QuanTest is set to 5. In Figure 11, we plot the accuracy, precision, recall, and F1 score of the CCQC model as a function of sampling cost. It is clear from this figure that the quality indicators of the QNN model are much more susceptible to the influence of QuanTest when the number of samples is small. As the sampling cost increases, the quality indicators rapidly rise in a similar trend, and when the number of samples exceeds $10^{4}$, their values stabilize close to the ideal expectations. This provides us with a reference for balancing physical resources and accurate testing. Considering the results in Figure 10, setting the number of samples to around $10^{4}$ during testing might be a relatively balanced choice in this case.
Answer to RQ4: The test samples generated by QuanTest can increase the sampling cost of the tested QNN system. When the number of samples is small, QuanTest may generate more test samples than expected, and the quality indicators of the QNN model are more susceptible to the influence of QuanTest.
Figure 11. Effect of QuanTest on the quality indicators of the CCQC model under different sampling costs.
[…]
5.2. Comparison with Black-box Baselines
The quantum-adapted adversarial attack strategies used by QuanTest have already been empirically evaluated in the classical world independent of entanglement guidance. To provide more empirical evidence of effectiveness, we compared QuanTest with black-box adversarial attacks on the binary classification task of Fashion-MNIST, following the same experimental setup as in RQ1. We select the state-of-the-art black-box adversarial attack method SimBA as the black-box baseline. Since it is difficult to achieve satisfactory results on the QNN system by directly using the default parameters, we use $\epsilon=0.1$ as the fixed step size of SimBA to compare more fairly with QuanTest. It should be noted that there is a significant difference between quantum states and classical images. Modifying quantum states as finely as manipulating classical images in pixel space is difficult. Therefore, SimBA here still processes classical data, without quantum-based adjustments.
Table 5 shows the comparison between QuanTest and the black-box baseline. It can be seen that although the black box baseline performs better than the random noise, there is still a significant gap compared to QuanTest. Unlike QuanTest, the black-box baseline finds it more difficult to generate test samples on the Fashion dataset, possibly because black-box attacks directly process classical images, leading to performance differences between different datasets. It is worth emphasizing that the QNN model can classify classical and quantum data (Biamonte et al., 2017). Adding perturbations to classical data is equivalent to adding perturbations to the corresponding quantum states (Lu et al., 2020), so QuanTest only needs to consider the usage scenarios for quantum data. End-to-end black-box adversarial attacks aimed at classical data cannot be directly applied to quantum data. Therefore, the usage scenarios of QuanTest and end-to-end black-box attacks are different.
Table 5. Comparison of QuanTest and black-box baseline on test sample generation rate and average trace distance.
Data Model Gen_Rate(%) ATD
SimBA QuanTest (DLFuzz-based) SimBA QuanTest (DLFuzz-based)
Fashion[1,4] CCQC 7.43 87.79 0.3136 0.3428
QCL 7.57 96.48 0.3224 0.3327
QCNN 13.71 92.73 0.3370 0.2947
MNIST[0,1,2] CCQC 16.23 70.95 0.3261 0.2852
QCL 19.11 71.52 0.3313 0.2467
QCNN 24.51 79.24 0.3294 0.1984
MNIST[3,6] CCQC 13.54 88.44 0.3588 0.3117
QCL 21.86 91.70 0.3521 0.2811
QCNN 20.68 96.87 0.3462 0.2881
Answer to RQ5: QuanTest has a better performance than the black box baseline. Note that QuanTest and end-to-end black box attacks have different usage scenarios.
5.3. Selection of Random Baseline
As existing work on testing QNN systems is still blank, we selected random coherent noise as the baseline to evaluate QuanTest in RQ1. Considering that random noise of smaller strength per iteration may have a higher test sample generation rate, we explore the generation of test samples using random coherent noise at different $\sigma$ settings on the Fashion dataset, following the same experimental settings in RQ1. The results are shown in Table 6.
We find that smaller values of $\sigma$ indeed lead to a higher generation rate of random noise, but this still has a significant gap compared to QuanTest’s performance. However, it is imperative to note that the decrease in strength means an increase in the number of iterations, which will result in greater time overhead. In fact, within the same time cost as QuanTest, more test samples are generated by random noise with $\sigma=0.02$ than other strength settings. Additionally, the time required for random noise with $\sigma=0.02$ to reach the same similarity limit as QuanTest is also closer. So it is a fairer setting to select the random noise of $\sigma=0.02$ as the baseline.
Table 6. Comparison of random baselines with different noise strengths on test sample generation rate and average trace distance.
Model Gen_Rate(%) ATD
$\sigma=0.02$ $\sigma=0.01$ $\sigma=0.005$ $\sigma=0.02$ $\sigma=0.01$ $\sigma=0.005$
CCQC 0.62 0.72 1.52 0.2143 0.2517 0.2727
QCL 0.83 1.18 1.68 0.3138 0.2733 0.2503
QCNN 2.49 2.13 2.32 0.2669 0.2696 0.2713
5.4. Scalability
The adversarial attack module used in QuanTest is configurable, which means that QuanTest can scale various adversarial attack strategies to generate diverse test samples without the need for quantum adaptation for each strategy. However, this scalability may be limited in end-to-end black-box attack strategies, as effective black-box attack methods specifically targeting quantum states have not yet been proposed. Additionally, based on our entire evaluation, QuanTest also has the ability to scale to high-dimensional quantum systems (i.e., larger networks and datasets) and a greater number of classes. Therefore, QuanTest’s testing targets are not limited to QNN systems designed for current NISQ. In future work, QuanTest will attempt to extend testing tasks to a broader range of architectures. It is worth noting our proposed QEA metric, as it relies solely on entanglement measure, theoretically QEA can be extended to other structural coverage test objectives for QNN systems.